OpenStudy (anonymous):

Quadratic Equations x^2-3x+2=0

OpenStudy (anonymous):

x^2 -2x -x + 2 = 0 (x - 2)(x-1) = 0

OpenStudy (anonymous):

x = 2,1

OpenStudy (anonymous):

yeah ishaan hi how u doin

OpenStudy (anonymous):

hey sheg, I am doing good what about you??

OpenStudy (anonymous):

m also doin good

OpenStudy (anonymous):

:) Nice

OpenStudy (anonymous):

but it was an easy question.